11. A 100 BHP diesel engine has a brake thermal efficiency of 30% and jacket water loss of 34%. The engine is to be cooled by means of a cooling tower having a range of 27oC. Find the mass of water leaving the tower.


1
Expert's answer
2022-03-13T15:15:38-0400

"Q_{Out} = \eta \times Q_{in}=0.3\times 74 .57 = 22.37 kW"


 mass of water leaving the tower.


"Q_{out}= m\times C_p\times \triangle T"


"m = \frac{22370}{1\times300}=74.56kg"